Solve the inequality. express the answer using interval notation.
3 * |(1/2)x + 1| + 2 is less then or equal to 47

Answer by stanbon(75887)

Solve the inequality. express the answer using interval notation.
3 * |(1/2)x + 1| + 2 is less then or equal to 47
------------------------
3 * |(1/2)x + 1| + 2 <= 47
---------------
3 * |(1/2)x + 1| <= 45
--------------
|(1/2)x + 1| <= 15
-15 <= (1/2)x+1 <= 15
-16 <= (1/2)x <= 14
----
-32 <= x <= 28
=================
Cheers,
Stan H.
=================